an open rectangle box with a length of 12 and a width of 9. the corners are cut out of the rectangle and the flaps are used to make an open, square box.
list the length as a function of x. f(x)=.....
list the width as a function of x. f(x)=....
list the equation for the volume as v(x)=....

Answer by ankor@dixie-net.com(22740)

an open rectangle box with a length of 12 and a width of 9. the corners are cut
out of the rectangle and the flaps are used to make an open, square box.
:
Let x = side of the corner cut-out
:
list the length as a function of x. f(x)=..L...
L = 12 - 2x
:
list the width as a function of x. f(x)=..W..
W = 9 - 2x
:
list the equation for the volume as v(x)=..V..
v = x(12-2x)(9-2x)
v = x(108 - 42x + 4x^2)
v = 4x^3 - 42x^2 + 108x